OpenStudy (anonymous):

find the following ^4sqrt2401

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt[4]{2401}\] \[\Large \sqrt[4]{7^4}\] \[\Large 7\] So \[\Large \sqrt[4]{2401}=7\]